Integrate fc2x (2x + 4)ex2+4x–3 dx
Answer #1

Guvern fons elles da put x +42-3 = n. differentiate BS: or za dat 4 dx = du (2x+y)dx = du. dr. du (2x+4) u -> 4) е du povers